OpenStudy (anonymous):

0 -5 3 -2 If A=(-2 -1 5 9) and B=﹛4 0﹜ -6 1 , What is the product of AB?

OpenStudy (anonymous):

0 -5 B= 3 -2 4 0 -6 1

OpenStudy (anonymous):

It is a matrix.

OpenStudy (helder_edwin):

\[\large \begin{pmatrix} -2 & -1 & 5 & 9 \end{pmatrix}\cdot \begin{pmatrix} 0 & -5\\ 3 & -2\\ 4 & 0\\ -6 & 1 \end{pmatrix}= \begin{pmatrix} -37 & 21 \end{pmatrix} \]

OpenStudy (anonymous):

Thank you @helder_edwin! Can you please tell me how did you calculate it?

OpenStudy (helder_edwin):

the first matrix if 1by 4 and the second is 4 by 2 so the product is 1 by 2. you multiply each row of the first matrix with each column of the second matrix. so, for example, \[\large (\color{red}{-2})\times\color{blue}{0}+(\color{red}{-1})\times\color{blue}{3}+ \color{red}{5}\times\color{blue}{4}+\color{red}{9}\times(\color{blue}{-6})=-37 \]

terenzreignz (terenzreignz):

Matrix multiplication gives me headaches :> Knowing dot-products makes things sorta simpler to understand...